### v1

# On unsolvable equations of prime degree

Abstract.
Kronecker observed that either all roots or only one root of a solvable irreducible equation of odd prime degree with integer coefficients are real. This gives a possibility to construct specific examples of equations not solvable by radicals. A relatively elementary proof without using the full power of Galois theory is due to Weber. We give a rather short proof of Kronecker’s theorem with a slightly different argument from Weber’s. Several modern presentations of Weber’s proof contain inaccuracies, which can be traced back to an error in the original proof. We discuss this error and how it can be corrected.
Key words and phrases:
Solvability by radicals, irredicible polynomials of prime degree, Galois theory, Kronecker, Weber
1991 Mathematics Subject Classification:
Primary 12F10, Secondary 12-03 01A55
Introduction
One of the main objectives of any course in Galois theory is a proof of the existence of polynomials which can not be solved by radicals, showing that similar formulas as for quadratic, cubic or quartic equations can not exist for higher degree equations. Today this kind of problems mainly has a pedagogical and historical value. Galois theory is an extremely important part of mathematics and therefore one of the fundamental parts of mathematical education. Solvability of polynomial equations is used as an illustration of the effectivity of abstract algebraic methods in solving concrete, important and interesting mathematical problems. It is usually presented using the full power of Galois theory, that is, the Galois correspondence and the relations between solvable equations and solvable groups. A hint that a more elementary approach is possible, is given by the circumstance that the original impossibility proof of Abel predates the work of Galois.
Specific examples of unsolvable equations are easily obtained from the fact that for an irreducible equation of odd prime degree with integer coefficients all of its roots or only one are real, an observation made by Kronecker [K2]. A relatively elementary proof, using rather limited knowledge related to the field extensions, was given by Weber [WW1].
Weber’s proof occurs in the algebra text book [Nag] by Trygve Nagell. Thanks to this book in Swedish the first Author became aware of the possibility to prove unsolvability of the quintic in an “elementary” way and transformed it into exercise 13.6 in [Brz]. Unfortunately, the included solution is incorrect. We were led to a closer study of Nagell’s proof and to a search for its origin. We found several recent presentations of Weber’s proof in text books [Pra, MJP], which are based on the English translation [DA] of the popular account by Dörrie [Dör]. Several objections to [DA] and expositions following it have been raised, see [PC, Sko]. The ultimate source for the inaccuracies observed is an error in Weber’s proof. The problem lies in reducible radicals (radicals of the form $\alpha=\sqrt[q]{a}$ with $X^{q}-a$ reducible). The need to deal with them also complicates the variant of the proof we developed.
In this paper we first place Kronecker’s theorem in a historical context. We discuss Kronecker’s arguments for the theorem and trace the history of “elementary” proofs. In the second section we give our short proof of the theorem. In the final section we discuss the error in Weber’s proof, and ways in which the proof can be corrected.
1. History
After Cardano published the solution of cubic and quartic equations in 1545, attempts were made to find solutions of quintic equations, expressing their roots in a similar way as function of the coefficients, by formulas involving only arithmetical operations and radicals. The work of Waring, Vandermonde and especially Lagrange round 1770 suggested that such formulas for equations of degree 5 probably do not exist. This was proven to be so in the work of Paulo Ruffini and Niels Henrik Abel between 1799 and 1826, see the expository paper by Rosen [Ros].
Much of the effort of Abel, Galois and the mathematicians continuing their work, was concentrated on the characterization of the equations solvable by radicals. The problem was clearly formulated by Abel in his unfinished memoir “Sur la resolution algébrique des équations” [Ab]. He also indicated the way to arrive at its solution. According to Abel one should find all equations of a given degree which are algebraically solvable. In the course of his investigations he arrives at “several general propositions about the solvability of equations and about the form of their roots” [Ab, p. 219].
The investigation of the form of the roots of a solvable equation was taken up by Kronecker, see the interesting paper by Petri and Schappacher [PS]. Three years after his first note on the subject [K1], Kronecker observes in [K2] the following:
Wenn eine irreductible Gleichung mit ganzzahligen Coëfficienten auflösbar und der Grad derselben eine ungrade Primzahl ist, so sind entweder alle ihre Wurzeln oder nur eine reell.  [^1] If an irreducible equation with integer coefficients is solvable and the degree of it is an odd prime, then either all of its roots or only one are real.
This is an easy to formulate special case of the general result:
Wenn eine Gleichung — deren Grad eine ungrade Primzahl $\mu$ ist, deren Coëfficienten rationale Functionen irgend welcher reeller Größen $A,B,C,\dots$ also selbst reell sind und welche endlich nicht in Factoren niederen Grades zerlegt werden kann, so daß deren Coëfficienten wiederum rationale Functionen von $A,B,C,\dots$ wären — durch eine explicite algebraische Function jener Größen $A,B,C,\dots$ erfüllt wird, so sind entweder alle ihre Wurzeln , oder nur eine derselben reell.  [^2] If an equation — whose degree is an odd prime $\mu$, whose coefficients are rational functions of any real quantities $A,B,C,\dots$ and therefore themselves real and which finally cannot be decomposed in factors of lower degree so that their coefficients are again rational functions of $A,B,C,\dots$ — is satisfied by an explicit algebraic function of those quantities $A,B,C,\dots$ , then either all of its roots, or only one of them are real.
The modern, much shorter way to express this statement is that an irreducible polynomial of (odd) prime degree over a real number field has exactly one real root or only real roots. In his later work Kronecker used the term domain of rationality. In his influential algebra book [We] Weber follows Dedekind and writes field, but adds that the term domain of rationality can be useful to express that in a certain problem the elements of the field should be considered as known or rational.
As Kronecker clearly states, he found his results from the study of the form of the roots of solvable polynomials. This study was continued by Weber [We], Wiman [Wi] and recently by Edwards [Ed]. In the penultimate section of the first volume of [We] Weber connects the number of real roots to the number of real roots of an auxiliary equation of degree $p-1$, and this seems to establish the Kronecker’s theorem, but Weber starts out from the result, referring to a simple proof earlier in his book. This simple proof was already mentioned by Kronecker himself [K2]:
Ich bemerke ferner, daß die angegebene Eigenschaft der irreductibeln auflösbaren Gleichungen $\mu$ten Grades nicht bloß aus der allgemeinen Form ihrer Wurzeln hervorgeht, sondern auch aus dem schon von Galois herrührenden Satze daß jede Wurzel einer solchen Gleichung sich als rationale Function von irgend zwei andern darstellen läßt“. Wenn nämlich diese Function nur reelle Coefficienten enthält, so folgt hieraus unmittelbar, daß alle Wurzeln reell sein müssen, sobald nur zwei derselben reell sind.  [^3] Furthermore I remark, that the stated property of irreducible solvable equations of degree $\mu$ does not only follow from the general shape of their roots, but also from the Theorem due to Galois, “that every root of such an equation can be represented as a rational function of any two others”. Namely, if this function only contains real coefficients, then it follows immediately that all roots must be real as soon as only two of them are real.
Galois’ result is in modern formulation:
Theorem 1 (Galois’ Theorem).
An irreducible polynomial equation of prime degree $p$ over a number field is solvable by radicals if and only if its splitting field is generated by any two of its zeroes.
For a proof see [Brz, Ex. 13.9]. Galois proved his result in the famous memoir that was rejected in 1831 by the Académie des Sciences. Poisson complained in the report that it did not contain, as the title of the Memoir promised, the condition of solvability of equations by radicals. To decide whether a given equation of prime degree is solvable, one should first have to determine if this equation is irreducible, and then if one of its roots can be expressed as a rational function of two others. The condition, if it exists, should be verifiable by inspecting the coefficients of a given equation, or, at most, by solving other equations of lower degree (see [Ta, p. 120]).
Kronecker’s theorem gives a possibility to construct equations, which are not solvable by radicals: any irreducible polynomial equation of odd prime degree with integral coefficients having at least three real roots and also complex conjugate roots will do. One of the simplest examples is the polynomial $X^{5}-4X-2$, see Figure 1. It is irreducible by the Schönemann–Eisenstein criterion, published by Schönemann in 1846 and Eisenstein in 1850 (see [Cox]); the criterion was of course not known to Poisson.
1.51.5-1.5- 1.511-1- 10.50.5-0.5- 0.50.50.50.50.511111.51.51.51.544-4- 422-2- 22222
Figure 1. The graph of $f(X)=X^{5}-4X-2$
A similar example is given in Ian Stewart’s text book Galois Theory [Ste], but on the basis of Weber’s theorem [We, §186], that an irreducible polynomial $f(X)\in K[X]$, $K$ a real number field, of prime degree $p$ with $p-2$ real and 2 complex nonreal roots has the symmetric group $S_{p}$ as Galois group, see [Brz, Ex. 13.4]. We note that Galois’ theorem as stated above does not occur in [Ste].
As Kronecker [K2] already remarked, the two cases, only one or all roots real, are for $p\equiv 3\pmod{4}$ distinguished by the sign of the discriminant of the polynomial (the square of the product of all differences between roots). This follows from the fact that the discriminant of a polynomial with real coefficients (without multiple factors) is positive if the number of pairs of complex conjugate roots is even and negative if this number is odd. It also follows that for $p\equiv 1\pmod{4}$ only irreducible equations with positive discriminant can be solvable.
A proof of Kronecker’s theorem was given by Weber in the first volume of Weber–Wellstein’s Encyklopädie der Elementar-Mathematik [WW1, § 101], originally for the case $p=5$, and in the later editions for general $p$. Elementary mathematics here means roughly the mathematics taught in secondary school, and this “handbook for teachers and students” is mainly directed at teachers. A fourth, heavily revised edition [WW2] was prepared by Paul Epstein. He made almost no changes in the chapter on solvable equations, but modernized the terminology.
Weber’s proof is suitable for a university course in algebra which treats the solution of equations, but not Galois theory. Such a course was given since 1931 at Uppsala University by Trygve Nagell. In his textbook [Nag] Nagell treats the case $p=5$ following the first edition of Weber–Wellstein.
The insolvability of the quintic was included as one of the hundred problems in Heinrich Dörrie’s book Triumph der Mathematik [Dör], intended for a general audience. It contains a variety of problems, like Kirkman’s schoolgirl problem, or the length of the polar night, but also the transcendence of $\pi$. Dörrie was one of the first Ph.D students of Hilbert with a dissertation on the quadratic reciprocity law in quadratic number fields with class number one. Dörrie’s book has been translated into Japanese, Hungarian and English. The English translation [DA] is unfortunately marred by strange terminology: Körper is not translated as field but as group, Adjunktion as substitution. For the insolvability of equations Dörrie follows the 1922 edition of Weber–Wellstein [WW2]. His version is also the basis for the expositions in [PC], [MJP] and [Pra]. The last text inspired the proof in [Sko]. None of the sources mentioned cites Weber.
2. A proof of Kronecker’s Theorem
We suppose that all fields in this section have characteristic zero. As a reference to some elementary results in Galois theory we use [Brz].
We recall that a field extension $K\subseteq L$ is radical if there is a chain
(1) $K=K_{0}\subseteq K_{1}\subseteq\cdots K_{i-1}\subseteq K_{i}\subseteq\cdots\subseteq K_{n}=L$
[…]
We stress the fact that in the definition of a radical extension we do not assume that the polynomials $X^{q_{i}}-\alpha_{i}^{q_{i}}$ are irreducible over $K_{i-1}$. If $X^{q}-\alpha^{q}$ is irreducible, then $\alpha=\sqrt[q]{a}$ is called an irreducible radical. Otherwise we call it reducible.
Let $q$ be a prime number and $K$ a number field. The splitting field of $X^{q}-a\in K[X]$ over $K$ is $K^{\prime}=K(\alpha,\varepsilon)$, where $\alpha\in K^{\prime}$ is a zero of $X^{q}-a$ and $\varepsilon$ a primitive $q$-th root of unity [Brz, Ex. 5.11]. All zeroes of $X^{q}-a$ are $\varepsilon^{i}\alpha$ for $i=0,\dots,q-1$. If no zero lies in $K$, or in other words, if $a$ is not the $q$-th power of an element in $K$, then $X^{q}-a$ is irreducible over $K$ [Brz, Ex. 5.12] and $[K(\alpha):K]=q$. If $X^{q}-a$ has a root $\alpha\in K$, then by the substitution $X=\alpha Y$ the equation becomes $Y^{q}-1=0$. Therefore $K^{\prime}=K(\varepsilon)$ with $\varepsilon$ a primitive $q$-th root of unity. Notice that the minimal polynomial $X^{q-1}+\dots+X+1$ of $\varepsilon$ over $\mathbb{Q}$ may be reducible over $K_{i-1}$. By the Theorem on natural irrationalities [Brz, Ex. 9.22] the extension $K^{\prime}\subseteq K$ is Galois and its group $G(K^{\prime}/K)$ is a subgroup of some order $d$ of the cyclic Galois group $G(\mathbb{Q}(\varepsilon)/\mathbb{Q})$ of order $q-1$. Therefore $[K^{\prime}:K]$ is equal to $d$, a divisor of $q-1$. In particular, if $\varepsilon\in K$, then $d=1$ and $K^{\prime}=K$. If $K^{\prime}\neq K$ and has degree $d<q$ over $K$, then the extension is normal and cyclic.
If $f(X)\in K[X]$ is solvable by radicals, and if $K\subseteq K^{\prime}$ is an extension, then $f(X)\in K^{\prime}[X]$ is also solvable by radicals. Indeed, we obtain a radical extension $K^{\prime}\subseteq L^{\prime}$ if we starting from the chain (1) define $K^{\prime}_{0}=K^{\prime}$ and $K^{\prime}_{i}=K^{\prime}_{i-1}(\alpha_{i})$, $i=1,\dots,n$ and remove unnecessary fields from $K^{\prime}=K^{\prime}_{0}\subseteq K^{\prime}_{1}\subseteq\cdots\subseteq K^{\prime}_{n}=L^{\prime}$.
We use the following lemmas.
Lemma 2 (Nagell’s Lemma, see [Brz, Ex. 4.13]).
If $f\in K[X]$ has prime degree $p$ and is irreducible over $K$, but reducible over a field extension $K\subset L$, then $p\mid[L:K]$.
Remark 3.
The Lemma is formulated differently in [Nag] and [Brz], but the proofs given establish the present result.
Lemma 4 ([Brz, Ex. 7.8]).
Let $K\subset L$ be a normal extension and let $f(X)$ be a monic polynomial irreducible over $K$ but reducible over $L$. Then all irreducible factors of $f(X)$ in $L[X]$ have the same degree.
In the following we are interested in solvability of polynomials with real coefficients. We say that a number field $K\subseteq\mathbb{C}$ is conjugation invariant if complex conjugation is an automorphism of this field. Notice that if $K$ is a conjugation invariant field, then its extension $K(\alpha)$, where $\alpha\in\mathbb{C}$, is conjugation invariant if and only if $\bar{\alpha}\in K(\alpha)$ (see [Brz, T.4.2]).
Lemma 5.
Let $K\subseteq L$ be a radical extension of a conjugation invariant field. Then there exist a chain $K=K_{0}\subseteq K_{1}\subseteq\cdots\subseteq K_{m}$ of simple radical extensions with $L\subseteq K_{m}$, where all fields $K_{i}$ are conjugation invariant.
Proof.
We use induction with respect to the number $n$ of simple radical extensions between $K$ and $L$. Assume that the Lemma is true for every radical extension of $K$ by a chain of less than $n\geq 1$ of simple radical extensions. The base case, when there is no extension, is trivially true.
Let $L$ be a radical extension of $K$ by $n$ simple radical extensions. Let $L^{\prime}\subseteq L$ be the last extension. Then $K\subseteq L^{\prime}(\alpha^{\prime})=L$, where $\alpha^{\prime q}=a^{\prime}\in L^{\prime}$ and $L^{\prime}$ is a radical extension of $K$ by (at most) $n-1$ simple radical extensions. By the induction hypothesis we have $L^{\prime}\subseteq K^{\prime}_{m^{\prime}}$, where $K=K^{\prime}_{0}\subseteq K^{\prime}_{1}\subseteq\cdots\subseteq K^{\prime}_{m^{\prime}}$ is a chain of simple radical extensions with all fields $K^{\prime}_{i}$, $i=1,\ldots,m^{\prime}$, conjugation invariant.
Define $\rho^{\prime}=\alpha^{\prime}\bar{\alpha}^{\prime}$. Then $\rho^{\prime q}=\alpha^{\prime q}(\bar{\alpha}^{\prime})^{q}=a^{\prime}\bar{a}^{\prime}\in K^{\prime}_{m^{\prime}}$. As $\rho^{\prime}$ is real, the fields $K_{m^{\prime}+1}=K_{m^{\prime}}(\rho^{\prime})$ and $K^{\prime}_{m^{\prime}+2}=K^{\prime}_{m^{\prime}+1}(\alpha^{\prime})=K^{\prime}_{m^{\prime}}(\rho^{\prime},\alpha^{\prime})=K^{\prime}_{m^{\prime}}(\alpha^{\prime},\bar{\alpha}^{\prime})$ are conjugation invariant. Of course, we have $L=L^{\prime}(\alpha^{\prime})\subseteq K^{\prime}_{m^{\prime}+2}=K^{\prime}_{m}(\rho^{\prime},\alpha^{\prime})$. In the chain
$K=K^{\prime}_{0}\subseteq\cdots\subseteq K^{\prime}_{m^{\prime}}\subseteq K^{\prime}_{m^{\prime}+1}\subseteq K^{\prime}_{m^{\prime}+2}=K^{\prime}_{m^{\prime}}(\rho^{\prime},\alpha^{\prime})$
all fields are conjugation invariant and $L\subseteq K^{\prime}_{m^{\prime}+2}$. This proves the induction step. ∎
Theorem 6 (Kronecker’s Theorem).
Let $K$ be a conjugation invariant field. Suppose that an irreducible polynomial equation $f(X)\in K[X]$ of odd prime degree $p$ with real coefficients is solvable by radicals. Then exactly one root or all roots of the polynomial are real.
Proof.
By Lemma 5 we may assume that there is a conjugation invariant chain $K=K_{0}\subseteq K_{1}\subseteq\cdots\subseteq K_{n}=L$ of simple radical extensions of prime degree with $K_{f}\subseteq L$. There exist an $i$ such that $f(X)$ is irreducible in $K_{i-1}$ but reducible in $K_{i}=K_{i-1}(\alpha)$, where $\alpha^{q}=a\in K_{i-1}$. By Nagell’s Lemma 2 the degree $p$ of $f(X)$ divides $d=[K_{i}:K_{i-1}]$.
If $X^{q}-a$ is irreducible over $K_{i-1}$, then $q=p$. In this case we may assume that $K_{i-1}$ contains a primitive $p$-th root of unity $\varepsilon_{p}$, for otherwise we first adjoin $\varepsilon_{p}$, while the polynomial $f(X)$ remains irreducible, again by Nagell’s Lemma, as the minimal polynomial of $\varepsilon_{p}$ over $K_{i-1}$ divides $X^{p-1}+X^{p-2}+\cdots+1$ Under the assumption the extension $K_{i-1}\subset K_{i}$ is normal. If $X^{q}-a$ is reducible, then $K_{i}=K_{i-1}(\varepsilon_{q})$ is a normal extension of $K_{i-1}$.
Since the extension $K_{i-1}\subset K_{i}$ making $f(X)$ reducible is normal, the degrees of all irreducible factors of $f(X)$ are equal by Lemma 4, so the factors are linear and $f(X)$ has all its zeroes $\beta_{1},\dots,\beta_{p}$ in $K_{i}$. If $[K_{i}:K_{i-1}]=p$, then $K_{i}=K_{i-1}(\beta_{k})$ for all $k$. If $[K_{i}:K_{i-1}]=d>p$, then the Galois group $G(K_{i}/K_{i-1}(\beta_{k}))$ has order $\frac{d}{p}$ and this group is the unique subgroup of this order of the cyclic group $G(K_{i}/K_{i-1})$. Therefore all the fields $K_{i-1}(\beta_{k})$, $k=1,\dots,p$, are equal as they are the field fixed by this subgroup (see [Brz, T.9.1]). So in all cases the splitting field $(K_{i-1})_{f}=K_{i-1}(\beta_{1},\dots,\beta_{p})$ of $f(X)$ over $K_{i-1}$ is equal to $K_{i-1}(\beta_{k})$, for all $k$.
Because $f(X)$ has real coefficients, it has a real root. Suppose that at least two roots, $\beta_{1}$ and $\beta_{2}$, are real. The Galois group $G((K_{i-1})_{f}/K_{i-1})$ is cyclic of order $p$. Every nontrivial automorphism is a permutation of the zeroes $\beta_{k}$ and must be a cycle of length $p$. Since the group is transitive, we may choose as generator a cycle $\sigma=(\beta_{1},\beta_{2},\dots,\beta_{p})$ starting with the real roots $\beta_{1},\beta_{2}$. Because the powers $\beta_{1}^{j}$ form a basis of $(K_{i-1})_{f}$ over $K_{i-1}$, we can write with uniquely determined coefficients $c_{j}\in K_{i-1}$
(2) $\beta_{2}=c_{0}+c_{1}\beta_{1}+\ldots+c_{p-1}\beta_{1}^{p-1}.$
Since $\beta_{1}$ and $\beta_{2}$ are real, taking complex conjugation gives,
$\beta_{2}=\bar{c}_{0}+\bar{c}_{1}\beta_{1}+\ldots+\bar{c}_{p-1}\beta_{1}^{p-1},$
so $\bar{c}_{i}=c_{i}$ for $i=0,1,\ldots,p-1$. Thus all the coefficients $c_{i}$ are real.
Applying the automorphism $\sigma$ to equation (2), we get
$\beta_{3}=c_{0}+c_{1}\beta_{2}+\ldots+c_{p-1}\beta_{2}^{p-1},$
showing that $\beta_{3}$ is also real. By repeating this last argument we find that all roots of $f(X)$ are real, if more than one root is real. ∎
3. Weber’s proof
As noted in the Introduction, there is an error in Weber’s proof. The problem is caused by reducible radicals. Let in a (not necessarily conjugation invariant) chain of simple radical extensions the first radical making $f(X)$ reducible be $\alpha=\sqrt[q]{a_{i}}$, with $q$ prime. Weber (and Epstein) [WW1, WW2] as well as Nagell [Nag] claim that $q=p$, but this does not hold if $\alpha$ is a reducible radical. Epstein and Dörrie give the following variant of Nagell’s Lemma ([WW2, § 111, 5] and [Dör, § 24, Satz IV]).
Proposition 7.
An irreducible polynomial $f(X)$ of prime degree $p$ can only become reducible through adjunction of a root of an [irreducible] equation $\varphi(X)=0$, whose degree is divisible by $p$.
We write the word irreducible in brackets, because it occurs in [Dör], but it is left out in [WW2], although the proof in [WW2] uses the assumption that the equation $\varphi(X)=0$ is irreducible. The Proposition is applied to the equation $X^{q}=a_{i}$, with the conclusion that $q=p$. Irreducibility of this equation is not assumed, not even implicitly, as the next step is an argument that $X^{p}-a_{i}$ is irreducible.
Dörrie [Dör] apparently noticed the error. To be able to use the correctly stated Proposition 7 he requires the radicals in the chain to be irreducible. From the point of view of writing down radical expressions for the roots of an equation it is quite natural to require irreducibility in the definition of solvability. This gives the same class of polynomials: a solvable polynomial is also solvable by irreducible radicals. The proof hinges on Gauss’ result that the roots of unity are expressible by irreducible radicals. For a short proof we refer to [vdW, § 62] and for a longer exposition [Tsch, IV, § 3]; see also the discussion in [Ste]. In the first edition of Weber–Wellstein [WW1] Weber took this fact up in an Appendix. In later editions the Appendix was moved to the main text, with its conclusion that this gives the same class of solvable polynomials, and reducible radicals were explicitly allowed.
At an early stage in his proof Dörrie [Dör] makes the chain conjugation invariant:
Ferner wollen wir mit jedem adjungierten Radikal unsere Kette, das noch keine Zerlegung von $f(x)$ ermöglicht, auch gleich das komplex konjugierte Radikal adjungieren. Das ist vielleicht überflüssig, sicher aber nicht schädlich. [^4] Furthermore, with each adjoined radical of our chain, which not yet makes it possible to factorise $f(x)$, we will also adjoin at the same time the complex conjugate radical. That is maybe superfluous, but certainly not harmful.
But actually harm can be done, as the complex conjugate radical may be reducible.
Example 8.
[…]
$X^{5}+X^{4}-4X^{3}-3X^{2}+3X+1=0\;.$
If the base field is $\mathbb{Q}$, then $\zeta$ satisfies an irreducible equation of degree $10=2\cdot 5$. This illustrates the statement about the degree of $\varphi(X)$ in Proposition 7.
If the chain of simple radical extensions starts with $K_{1}=\mathbb{Q}(\alpha)$, where e.g. $\alpha=\zeta\sqrt[11]{37}$ is an irreducible radical , with $11]{37}$ the real root and $\zeta$ a primitive $11$-th root of unity, then the splitting field of $f(X)$ is contained in $K_{2}=\mathbb{Q}(\alpha,\bar{\alpha})$ and $\bar{\alpha}$ is a reducible radical over $K_{1}=\mathbb{Q}(\alpha)$ (satisfying $X^{11}-37=X^{11}-\alpha^{11}=0$).
The problem with reducible radicals was also commented on by Skopenkov [Sko, p. 183], who pointed out an error in the proof by Prasolov [Pra]. Prasolov follows Dörrie, but he does not include irreducibility in the definition of solvability by radicals; nevertheless, he states without justification that the equation $X^{q_{i}}-a_{i}=0$, defining the extension $K_{i-1}\subset K_{i}$, is irreducible.
Weber’s and Dörrie’s argument can be rescued if the adjunction of a reducible radical making the polynomial $f(X)$ reducible is replaced by a chain of adjunctions of irreducible radicals. A different solution is to first adjoin all $q$-th roots of unity, for every root exponent $q$ occurring in the chain, and this via irreducible radicals. This is the approach chosen by Skopenkov [Sko] and also by Pan and Chen in their preprint [PC], where they fill a gap they found in Dörrie’s proof. Interestingly, they give a similar example as above with the quintic of example 8, but to illustrate that $f(X)$ can become reducible after adjoining the complex conjugate root, that is by an extension which is not conjugation invariant. It is clear that Weber considers adjoining $\alpha_{i}$ and $\bar{\alpha}_{i}$ as one step, introduced rather late in his proof. Only in the case that $f(X)$ splits, he adjoins first $\rho_{i}$ and then $\alpha_{i}$. As Dörrie follows Weber’s argument, he also considers adjoining $\alpha_{i}$ and $\bar{\alpha}_{i}$ as one step, although his text suggests otherwise.
Weber’s original arguments can also be adapted to handle reducible radicals. The start of the proof of Kronecker’s Theorem in [WW2] is similar to our arguments above. The $p$-th roots of unity are assumed to be adjoined, and it is shown that $f(X)$ splits in linear factors. The chain is made conjugation invariant. To finish Weber’s proof it remains to show the following lemma.
Lemma 9.
Let $f(X)\in K(X)$ be an irreducible polynomial of odd prime degree $p$ with real coefficients lying in a conjugation invariant field $K$ and let $K\subset L=K(\alpha)=K(\alpha,\bar{\alpha})$ be a normal simple radical extension such that $f(X)$ splits in linear factors over $L$. Then exactly one root or all roots of $f(X)$ are real.
Proof.
We distinguish two cases, depending on whether $\alpha=\sqrt[q]{a}$ is an irreducible radical or not.
Case I. If $\alpha$ is irreducible, then $q=p$. As the polynomial $f(X)$ has real coefficients, it has at least one real root $\beta_{0}$. All roots $\beta_{0},\dots,\beta_{p-1}$ can be expressed as linear combinations of $1,\alpha,\dots,\alpha^{p-1}$. In particular we have for the real root $\beta_{0}$ that
(3) $\beta_{0}=c_{0}+c_{1}\alpha+\dots+c_{p-1}\alpha^{p-1}\;.$
If $\alpha^{p}=a\notin\mathbb{R}$, then we consider $\rho=\alpha\bar{\alpha}$ with $\rho^{p}=a\bar{a}$. We have $K\subseteq K(\rho)\subseteq K(\alpha,\bar{\alpha})=K(\alpha)$. Then either $\rho\in K$ or $K(\rho)=K(\alpha)$. In the second case we replace the radical $\alpha$ by $\rho$ and get the same field. In case $a\in\mathbb{R}$ the same argument goes through, but brings nothing new: we may suppose that $\alpha\in\mathbb{R}$, which gives $\rho=\alpha^{2}$ and $\alpha=a^{-1}\rho^{\frac{p+1}{2}}$. Therefore we are left with two subcases, that $\alpha$ is real or that $\rho=\alpha\bar{\alpha}\in K$.
Case Ia. If $\alpha\in\mathbb{R}$, then complex conjugation of equation (3) gives that
$\beta_{0}=\bar{c}_{0}+\bar{c}_{1}\alpha+\dots+\bar{c}_{p-1}\alpha^{p-1}\;,$
where the $\bar{c}_{i}$ also lie in $K$. By uniqueness of the representation $\bar{c}_{i}=c_{i}$, or in other words, all $c_{i}$ are real. The other roots of $f(X)$ are
$\beta_{j}=c_{0}+c_{1}\varepsilon^{j}\alpha+\dots+c_{p-1}\varepsilon^{j(p-1)}\alpha^{p-1}\;.$
As $\bar{\varepsilon}^{j}=\varepsilon^{p-j}$, we find that $\bar{\beta}_{j}=\beta_{p-j}$, so the other roots come in complex conjugate pairs and in this case $\beta_{0}$ is the only real root.
Case Ib. If $\rho=\alpha\bar{\alpha}\in K$ then $\bar{\alpha}=\frac{\rho}{\alpha}$ and $\bar{\alpha}^{i}=\frac{\rho^{i}}{\alpha^{i}}=\frac{\rho^{i}}{a}\alpha^{p-i}$, so in this case complex conjugation of equation (3) gives $c_{0}=\bar{c}_{0}$ and $\bar{c}_{i}=\frac{a}{\rho^{i}}c_{p-i}$. For the other roots we find $\bar{\varepsilon}^{ij}\bar{\alpha}^{i}=\frac{\rho^{i}}{a}\varepsilon^{(p-i)j}\alpha^{p-i}$, which implies $\bar{\beta}_{j}=\beta_{j}$. Therefore all roots are real. This explicit computation makes the argument similar to the first case. Weber argues that $\beta_{0}=\psi(\alpha)=\bar{\psi}(\bar{\alpha})=\bar{\psi}(\frac{\rho}{\alpha})$, where $\bar{\psi}$ is the rational function obtained from $\psi$ by replacing all coefficients by their complex conjugates, and that $\psi(\alpha)=\bar{\psi}(\frac{\rho}{\alpha})$ still holds when $\alpha$ is replaced by another root $\alpha_{j}$; as $\bar{\alpha}_{j}=\frac{\rho}{\alpha_{j}}$ for any root of $X^{p}-a$, the equality $\psi(\alpha_{j})=\bar{\psi}(\bar{\alpha}_{j})$ holds.
Case II. If $\alpha$ is a reducible radical, then $\alpha^{q}=a_{0}^{q}$ for some $a_{0}\in K$. Therefore $\alpha=\varepsilon_{q}a_{0}$ for some primitive $q$-th root of unity and $K(\alpha)=K(\varepsilon_{q})$. The degree $d=[K(\varepsilon_{q}):K]$ is a multiple of $p$. We can write
$\beta_{0}=c_{0}+c_{1}\varepsilon_{q}+\dots+c_{d-1}\varepsilon_{q}^{d}=\psi(\varepsilon_{q})\;.$
As in case Ib, we have $\beta_{0}=\psi(\varepsilon_{q})=\bar{\psi}(\bar{\varepsilon}_{q})=\bar{\psi}(\frac{1}{\varepsilon_{q}})$ and the equality $\psi(\varepsilon_{q})=\bar{\psi}(\frac{1}{\varepsilon_{q}})$ still holds when $\varepsilon_{q}$ is replaced by another root of the equation of degree $d$ satisfied by $\varepsilon_{q}$ , and therefore $\psi(\varepsilon_{q})=\bar{\psi}(\bar{\varepsilon}_{q})$ holds. We find that $\bar{\beta}_{j}=\beta_{j}$ for all roots; if $d=kp$, then $k$ different $q$-th roots of unity give this conclusion for the same root of $f(X)$. Also in this case all roots are real. ∎