### v2

# On unsolvable equations of prime degree

Abstract.
Kronecker observed that either all roots or only one root of a solvable irreducible equation of odd prime degree with integer coefficients are real. This gives a possibility to construct specific examples of equations not solvable by radicals. A relatively elementary proof without using the full power of Galois theory is due to Weber. We give a rather short proof of Kronecker’s Theorem with an argument that is slightly different from Weber’s. Several modern presentations of Weber’s proof contain inaccuracies, which can be traced back to an error in the original proof. We discuss this error and how it can be corrected.
Key words and phrases:
Solvability by radicals, irredicible polynomials of prime degree, Galois theory, Kronecker, Weber
1991 Mathematics Subject Classification:
Primary 12F10, Secondary 12-03 01A55
Introduction
One of the main objectives of any course in Galois theory is a proof of the existence of polynomials which can not be solved by radicals, showing that similar formulas as for quadratic, cubic or quartic equations can not exist for higher degree equations. Today this kind of problems mainly has a pedagogical and historical value. Galois theory is an extremely important part of mathematics and therefore one of the fundamental parts of mathematical education. Solvability of polynomial equations is used as an illustration of the effectivity of abstract algebraic methods in solving concrete, important and interesting mathematical problems. The fact that most equations cannot be solved by radicals, is usually presented using the full power of Galois theory, that is, the Galois correspondence and the relations between solvable equations and solvable groups. A hint that a more elementary approach is possible, is given by the circumstance that the original impossibility proof of Abel predates the work of Galois.
Specific examples of unsolvable equations are easily obtained from the fact that for an irreducible equation of odd prime degree with integer coefficients all of its roots or only one are real, an observation made by Kronecker [K2]. A relatively elementary proof, using rather limited knowledge related to the field extensions, was given by Weber [WW1].
Weber’s proof occurs in the algebra text book [Na] by Trygve Nagell. Thanks to this book in Swedish the first Author became aware of the possibility to prove unsolvability of the quintic in an “elementary” way and transformed it into exercise 13.6 in [Br]. Unfortunately, the included solution is incorrect. We were led to a closer study of Nagell’s proof and to a search for its origin. We found several recent presentations of Weber’s proof in text books [Pr, MJP], which are based on the English translation [DA] of the popular account by Dörrie []. Several objections to [DA] and expositions following it have been raised, see [PC, Sk]. The ultimate source for the inaccuracies observed is an error in Weber’s proof. The problem lies in reducible radicals (radicals of the form $\alpha=\sqrt[q]{a}$ with $X^{q}-a$ reducible). The need to deal with them also complicates the variant of the proof we developed.
In Section 1, we place Kronecker’s Theorem in a historical context. We discuss Kronecker’s arguments for the theorem and trace the history of “elementary” proofs. In Section 2, we discuss the criteria that make the proofs of Kronecker’s Theorem elementary. We also recall a number of algebraic results that, according to these criteria, are needed in its elementary proofs. Section 3 contains our proof of Kronecker’s Theorem. Finally, in Section 4, we discuss the mistake in Weber’s proof and in proofs inspired by his idea, as well as ways in which the proofs can be corrected.
1. History.
After Cardano published the solution of cubic and quartic equations in 1545, attempts were made to find solutions of quintic equations, expressing their roots in a similar way as function of the coefficients, by formulas involving only arithmetical operations and radicals. The work of Waring, Vandermonde and especially Lagrange round 1770 suggested that such formulas for equations of degree 5 probably do not exist. This was proven to be so in the work of Paulo Ruffini and Niels Henrik Abel between 1799 and 1826, see the expository paper by Rosen [Ro].
Much of the effort of Abel, Galois and the mathematicians continuing their work, was concentrated on the characterization of the equations solvable by radicals. The problem was clearly formulated by Abel in his unfinished memoir “Sur la resolution algébrique des équations” [Ab1]. He also indicated the way to arrive at its solution. According to Abel one should find all equations of a given degree that are algebraically solvable. In the course of his investigations he arrives at “several general propositions about the solvability of equations and about the form of their roots” [Ab1, p. 219].
The investigation of the form of the roots of a solvable equation was taken up by Kronecker, see the interesting paper by Petri and Schappacher [PS]. Three years after his first note on the subject [K1], Kronecker observes in [K2] the following:
Wenn eine irreductible Gleichung mit ganzzahligen Coëfficienten auflösbar und der Grad derselben eine ungrade Primzahl ist, so sind entweder alle ihre Wurzeln oder nur eine reell.  [^1] If an irreducible equation with integer coefficients is solvable and the degree of it is an odd prime, then either all of its roots or only one are real.
This is a special case of the general result:
Wenn eine Gleichung — deren Grad eine ungrade Primzahl $\mu$ ist, deren Coëfficienten rationale Functionen irgend welcher reeller Größen $A,B,C,\dots$ also selbst reell sind und welche endlich nicht in Factoren niederen Grades zerlegt werden kann, so daß deren Coëfficienten wiederum rationale Functionen von $A,B,C,\dots$ wären — durch eine explicite algebraische Function jener Größen $A,B,C,\dots$ erfüllt wird, so sind entweder alle ihre Wurzeln, oder nur eine derselben reell.  [^2] If an equation — whose degree is an odd prime $\mu$, whose coefficients are rational functions of any real quantities $A,B,C,\dots$ and therefore themselves real and which finally cannot be decomposed in factors of lower degree so that their coefficients are again rational functions of $A,B,C,\dots$ — is satisfied by an explicit algebraic function of those quantities $A,B,C,\dots$ , then either all of its roots, or only one of them are real.
At the time Kronecker wrote this, the concept of field was not yet defined. With it Kronecker’s result can be formulated in the following way.
Theorem 1 (Kronecker’s Theorem).
Let $K$ be a real field. Suppose that an irreducible polynomial $f(X)\in K[X]$ of odd prime degree $p$ is solvable by radicals. Then exactly one root or all roots of the polynomial are real.
As Kronecker clearly states, he found his results from the study of the form of the roots of solvable polynomials. This study was continued by Weber [We], Wiman [Wi] and recently by Edwards [Ed]. In the penultimate section of the first volume of his algebra text book [We] Weber connects the number of real roots to the number of real roots of an auxiliary equation of degree p1p-1, and this seems to establish the Kronecker’s Theorem, but Weber starts out from the result, referring to a simple proof earlier in his book. This simple proof was already mentioned by Kronecker himself [K2]:
[…]
Galois’ result can be formulated in modern language as follows.
Theorem 2 (Galois’ Theorem).
An irreducible polynomial equation of prime degree $p$ over a number field is solvable by radicals if and only if its splitting field is generated by any two of its roots.
Kronecker’s Theorem follows, just as Kronecker remarked, immediately, since a solvable irreducible polynomial of prime degree $p$ over a real field $K$ having at least two real roots $\alpha,\beta$ splits in the field $K(\alpha,\beta)$, so all its roots are real.
Galois proved this result in his famous “Mémoire sur les conditions de résolubilité des équations par radicaux” that was rejected in 1831 by the Académie des Sciences, and only published by Liouville in 1846, see [Ne]; for a proof see [Br, Ex. 13.9]. Poisson, the referee of the memoir, noted in his report (see [Ne, p. 146]) that an analogous proposition of Abel was printed after Abel’s death in Crelle’s Journal [Ab2]. The excerpt in question of Abel’s letter to Crelle from October 18, 1928 reads:
Si trois racines d’une équation quelconque irréductible, dont le degré est un nombre premier, sont liées entre elles de sorte que l’une de ces racines peut être exprimée rationnellement par les deux autres, l’équation en question sera toujours résoluble à l’aide de radicaux.  [^4] If three roots of an arbitrary equation of which the degree is a prime number are related to each other in such a way that any one of these roots may be expressed rationally by the other two, the equation in question will always be solvable with the help of radicals.
It gives only a sufficient condition. The report goes on to say that no proof has been published and that Galois’ proof of his necessary and sufficient condition is not satisfactory. Poisson complained that the memoir did not contain, as the title promised, the condition of solvability of equations by radicals. For, to decide whether a given equation of prime degree is solvable, one should first have to determine if this equation is irreducible, and then if one of its roots can be expressed as a rational function of two others. The condition, if it exists, should be verifiable by inspecting the coefficients of a given equation, or, at most, by solving other equations of lower degree (see [Ne, p. 148]).
Kronecker’s Theorem gives a possibility to construct equations that are not solvable by radicals: any irreducible polynomial equation of odd prime degree with integral coefficients having more than one real root and also complex conjugate roots will do. One of the simplest examples is the polynomial $X^{5}-4X-2$, see Figure 1. The function $f(X)$ has only two real extrema and the polynomial is irreducible by the Schönemann–Eisenstein criterion, published by Schönemann in 1846 and Eisenstein in 1850 (see [Co]); this criterion was of course unknown to Poisson.
1.5-1.51-10.5-0.50.50.5111.51.54-42-222
Figure 1. The graph of $f(X)=X^{5}-4X-2$
A similar example is given in Ian Stewart’s text book Galois Theory [St], but on the basis of Weber’s theorem [We, §186], that an irreducible polynomial $f(X)\in K[X]$, $K$ a real number field, of prime degree $p$ with p2p-2 real and 2 complex nonreal roots has as Galois group the symmetric group $S_{p}$, see [Br, Ex. 13.4]. We note that Galois’ theorem as stated above does not occur in [St], as in many other modern textbooks.
As Kronecker [K2] already remarked, the two cases, only one or all roots real, are for $p\equiv 3\pmod{4}$ distinguished by the sign of the discriminant of the polynomial (the square of the product of all differences between roots). This follows from the fact that the discriminant of a polynomial with real coefficients (without multiple factors) is positive if the number of pairs of complex conjugate roots is even and negative if this number is odd. It also follows that for $p\equiv 1\pmod{4}$ only irreducible equations with positive discriminant can be solvable.
A proof of Kronecker’s Theorem was given by Weber in the first volume of Weber–Wellstein’s Encyklopädie der Elementar-Mathematik [WW1, § 101], originally for the case $p=5$, and in the later editions for general $p$. Elementary mathematics here means roughly the mathematics taught in secondary school, and this “handbook for teachers and students” is mainly directed at teachers. A fourth, heavily revised edition [WW2] was prepared by Paul Epstein. He made almost no changes in the chapter on solvable equations, but modernized the terminology.
Weber’s proof is suitable for a university course in algebra that treats the solution of equations, but not Galois theory. Such a course was given since 1931 at Uppsala University by Trygve Nagell. In his textbook [Na] Nagell treats the case $p=5$ following the first edition of Weber–Wellstein.
The insolvability of the quintic was included as one of the hundred problems in Heinrich Dörrie’s book Triumph der Mathematik [], intended for a general audience. It contains a variety of problems, like Kirkman’s schoolgirl problem, or the length of the polar night, but also the transcendence of $\pi$. Dörrie was one of the first Ph.D students of Hilbert with a dissertation on the quadratic reciprocity law in quadratic number fields with class number one. Dörrie’s book has been translated into Japanese, Hungarian and English. The English translation [DA] is unfortunately marred by strange terminology: Körper is not translated as field but as group, Adjunktion as substitution. For the insolvability of equations Dörrie follows the 1922 edition of Weber–Wellstein [WW2]. His version is also the basis for the expositions in [PC], [MJP] and [Pr]. The last text inspired the proof in [Sk]. None of the sources mentioned cites Weber [^5] The latest version of the preprint [PC] contains a footnote stating that Weber earlier gave the same proof..
2. Preliminaries
2.1. What is an elementary proof?
Elementary certainly means avoiding deep general theory. In the context of Kronecker’s Theorem this excludes the fundamental theorem of Galois theory. What can be used depends on the target group of the exposition. Almost a century ago Dörrie wrote for a general audience. Consequently he started his chapter on Kronecker’s Theorem by explaining what a number field is and proved the facts about field extensions needed for Weber’s proof, but avoided the concept of group. The proof uses the theory of field extensions within the scope defined by the contents of a second course in algebra at a university level. For our notion of elementary we have students taking such a course in mind. The considerations can always be limited to algebraic number fields, that is, subfields $K$ of the complex numbers $\mathbb{C}$. Thus, it is possible to include field extensions by the roots of polynomials (consequently, their splitting fields), the description of simple extensions, and the behaviour of the degree according to the Tower Law. It is essentially no problem to discuss automorphism groups of fields, since it is simply one more example of groups, which are always discussed in such courses. Thus, also the notion of the Galois group of a polynomial $f(X)\in K[X]$ as the group of all automorphisms of its splitting field $K_{f}=K(\alpha_{1},\ldots,\alpha_{n})$, where $\alpha_{i}$ are all the roots of $f(X)=0$ in $\mathbb{C}$, may be accepted as elementary. This group we denote by $G(K_{f}/K)$. What is essential to avoid is the fundamental theorem of Galois theory (together with the Galois correspondence, in particular, between normal subgroups and normal extensions of the ground field), and as regards the notion of solvability, the correspondence between solvable groups and solvable extensions. But the notion of a solvable extension of fields is, of course, elementary. It is often presented as an example of the power of (abstract) algebraic language when one wants to formulate the problem of algebraic solvability and, in particular, when one wants to encourage the students to read a course in Galois theory in order to to find out why there is no formula for an algebraic solution of a general quintic.
2.2. Field extensions.
In this subsection, we recall several well-known results concerning this “elementary part” of the theory of (number) field extensions that are used in the discussion of elementary proofs of Kronecker’s Theorem. As a general reference for the notions and results that we mention we use [Br].
Let $K\subseteq L$ be a field extension. An element $\alpha\in L$ is algebraic if there is a nonzero polynomial $f(X)\in K[X]$ such that $f(\alpha)=0$. Zeros of univariate polynomials are called roots. Recall that a field extension $K\subseteq L$ is simple, if $L$ is obtained by adjoining one element $\alpha$, so $L$ is of the form $K(\alpha)$. This field is isomorphic to $K[X]/(p(X))$, with $p(X)$ the minimal polynomial of $\alpha$, that is, the monic polynomial of lowest degree, of which $\alpha$ is a root. Thus $K(\alpha)$ is a vector space over $K$, with basis $1,\alpha,\dots,\alpha^{n-1}$, where $n$ is the degree of the minimal polynomial $p(X)$. This is also the degree $K(\alpha):K]$ of the extension $K\subseteq K(\alpha)$. The previous statements are the contents of the Simple Extension Theorem [Br, T.4.2]. The Tower Law [Br, T.4.3] states that, if $K\subseteq L$ and $L\subseteq M$ are finite field extensions, then also $K\subseteq M$ is a finite field extension and $[M:K]=[M:L][L:K]$.
A splitting field $K_{f}$ for a polynomial $f(X)\in K[X]$ is a field $K_{f}$ that contains all roots $\alpha_{1},\dots,\alpha_{n}$ of $f$ and is generated over $K$ by these roots, that is, $K_{f}=K(\alpha_{1},\dots,\alpha_{n})$. Recall that a field extension $K\subseteq L$ is normal if every irreducible polynomial over $K$ that has one root in $L$ necessarily has all the remaining roots in $L$. This property of splitting fields is usually not included in “elementary” algebra courses, but a proof that splitting fields of polynomials over a field $K$ coincide with finite normal extensions $L$ of this field is really elementary and uses only the structure of simple extensions and the Tower Law together with basic properties of isomorphism of fields (see the proof of [Br, T.7.1] on [Br, p. 120]).
We mention two properties of isomorphisms of fields, which play an essential role in several arguments. The first is often proved in the context of a theorem usually attributed to Kronecker, stating that for any field and any non-constant polynomial, there exists an extension field containing a root of this polynomial. The second is less known, but its proof is closely related to the first one.
Lemma 3.
(a) Let $K$ be a (number) field and let $\alpha,\alpha^{\prime}$ be two roots of an irreducible polynomial $f(X)\in K[X]$. Then there is an isomorphism $\varphi:K(\alpha)\rightarrow K(\alpha^{\prime})$ such that $\varphi(\alpha)=\alpha^{\prime}$ and $\varphi$ is the identity on $K$.
(b) Let $K\subset L$ be a normal extension and let $\alpha,\alpha^{\prime}\in L$ be two roots of an irreducible polynomial $f(X)\in K[X]$. Then there is an automorphism $\varphi$ of $L$ over $K$ such that $\varphi(\alpha)=\alpha^{\prime}$.
For a proof of (a) we refer to [Br, T.5.1], and for a proof of (b) to [Br, Ex. 7.4 (a)]. Notice that (b) says that the automorphism group of a normal field extension $L$ of a field $K$ acts transitively on the roots of any irreducible polynomial in $K[X]$ that has its roots in $L$.
We are interested in algebraic solvability of polynomials, which means expressing the roots using arithmetical operations and radicals. The problem is best formulated in terms of fields. A simple radical extensions is an extension $K\subseteq K(\alpha)$, where $\alpha^{q}=a\in K$ for some positive integer $q$, that is, $\alpha$ is a root of the binomial polynomial $X^{q}-a$. A field extension $K\subseteq L$ is radical if there is a chain
(1) $K=K_{0}\subseteq K_{1}\subseteq\cdots K_{i-1}\subseteq K_{i}\subseteq\cdots\subseteq K_{n}=L$
of simple radical extensions $K_{i}=K_{i-1}(\alpha_{i})$. We may assume that the $q_{i}$ are prime numbers, by using the fact that $\sqrt[rs]{a}=\sqrt[r]{\sqrt[s]{a}}$. A polynomial $f(X)\in K[X]$ is solvable by radicals, if a splitting field of $f(X)$ is contained in a radical extension $K\subseteq L$ [Br, p. 78].
We stress the fact that in the definition of a simple radical extension we do not assume that the polynomial $X^{q}-\alpha^{q}$ is irreducible over $K$. If $X^{q}-\alpha^{q}$ is irreducible, then $\alpha=\sqrt[q]{a}$ is called an irreducible radical. Otherwise, that is, if $X^{q}-a$ is reducible over $K$, then any $\alpha=\sqrt[q]{a}$ is called a reducible radical. A polynomial $f(X)\in K[X]$ is solvable by irreducible radicals, if $K_{f}\subseteq L$, where $K\subseteq L$ is a radical extension such that in the chain of simple radical extensions every $X^{q_{i}}-\alpha_{i}^{q_{i}}$ is irreducible over $K_{i-1}$. From the (constructive) point of view of writing down radical expressions for the roots of an equation it is quite natural to require irreducibility in the definition of solvability. But this gives the same class of polynomials: a solvable polynomial is also solvable by irreducible radicals, see e.g. [St, Ch. 21].
In the following sections, we need a few properties of the splitting fields of the binomials $X^{q}-a$ over number fields $K$. We gather them in the following Lemma.
Lemma 4.
Let $K$ be an algebraic number field, $q$ a prime number and $a\in K\setminus\{0\}$.
(a) The splitting field of $X^{q}-a\in K[X]$ over $K$ is $K^{\prime}=K(\alpha,\varepsilon)$, where $\alpha\in K^{\prime}$ is a root of $X^{q}-a$ and $\varepsilon$ a primitive $q$-th root of unity.
(b) The binomial $X^{q}-a$ is reducible over $K$ if and only if it has a root in $K$.
(c) If $\varepsilon\in K$ and $X^{q}-a$ is irreducible over $K$, then $K^{\prime}=K(\alpha)$ and $K^{\prime}:K]=q$. All automorphisms of $K^{\prime}$ over $K$ are mappings $\varphi(\alpha)=\varepsilon^{r}\alpha$ where $r=0,1,\ldots,q-1$. Thus the Galois group is the cyclic group $G(K^{\prime}/K)=\mathbb{Z}_{q}$.
(d) If $\varepsilon\notin K$ and $X^{q}-a$ is reducible over $K$, then $K^{\prime}=K(\varepsilon)$ is the splitting field of the polynomial $X^{q}-1$. Every automorphism of $K^{\prime}$ maps $\varepsilon$ onto a power of $\varepsilon$.
We will not present detailed proofs, but we comment on the points above.
(a) The equation $X^{q}-a=0$ has $q$ different solutions $\varepsilon^{r}\sqrt[q]{a}$, where $\varepsilon\neq 1$ denotes a (primitive) $q$-th root of unity, $r=0,\ldots,q-1$ and $\alpha=\sqrt[q]{a}$ is any fixed solution to the equation $X^{q}-a=0$. Thus the splitting field of the binomial $X^{q}-a$ over $K$ is equal to $K^{\prime}=K(\alpha,\varepsilon)$ ([Br, Ex. 5.11 (c)]).
(b) This is a result of Abel. The fact that the polynomial $X^{q}-a$ is reducible over $K$ if and only if $a$ is a $q$-th power in $K$ is proved in [Br, Ex. 5.12 (a)].
(c) If $\varepsilon\in K$ and $X^{q}-a$ is irreducible, then according to (b), $\alpha\notin K$, since $a$ is not a $q$-th power in $K$ ($\alpha^{q}=a$). The extension $K^{\prime}=K(\alpha)$ of $K$ is simple, of degree $q$ by the Simple Extension Theorem. According to (a), the equation $X^{q}-a=0$ has $q$ different solutions $\varepsilon^{r}\alpha$ for $r=0,1,\dots,q-1$. Every automorphism of $K^{\prime}$ maps $\alpha$ on a root of the same polynomial. Thus, we have $q$ different automorphisms mapping $\alpha$ onto $\varepsilon^{r}\alpha$. The exponents $r$ form the group $\mathbb{Z}_{q}$ of all residues $r$ modulo $q$ (as $\varepsilon^{q}=1)$.
(d) If $X^{q}-a$ has a root $\alpha\in K$, then it follows from (a) that $K^{\prime}=K(\varepsilon)$, where as before $\varepsilon$ denotes a primitive $q$-th root of unity. Thus $K^{\prime}$ is a splitting field of the polynomial $X^{q}-1$, since all $q$-th roots of unity are powers of $\varepsilon$. The minimal polynomial of $\varepsilon$ over $K$ divides the minimal polynomial of $\varepsilon$ over $\mathbb{Q}$, which has degree q1q-1; it is $(X^{q}-1)/(X-1)$ ([Br, Ex. 4.3 (f)], for the history of this result see [Co]). Every automorphism maps $\varepsilon$ onto another $q$-th root of unity, which is a power $\varepsilon^{k}$ for some $1\leq k\leq q-1$.
2.3. Auxiliary results.
The proofs of the following results, which are needed for Weber’s and our proof of Kronecker’s Theorem, do not depend on Galois theory.
Lemma 5 (Nagell’s Lemma).
If $f\in K[X]$ has prime degree $p$ and is irreducible over $K$, but reducible over a field extension $K\subset L$, then $p\mid[L:K]$.
The Lemma is formulated differently in [Na] and [Br, Ex. 4.13], but the proofs given there establish the present result. Weber [WW2] and Dörrie [] use Proposition 11 below with the same effect. The proof of the following Lemma can be found in [Br, Ex. 7.8]. It only depends on Lemma 3 (a) and the Tower Law.
Lemma 6.
Let $K\subset L$ be a normal extension and let $f(X)$ be a monic polynomial irreducible over $K$ but reducible over $L$. Then all irreducible factors of $f(X)$ in $L[X]$ have the same degree.
Kronecker’s Theorem is about solvability of polynomials with real coefficients. We need terminology and results adapted to this situation. We say that a number field $K\subseteq\mathbb{C}$ is conjugation invariant if complex conjugation is an automorphism of this field. Notice that if $K$ is a conjugation invariant field, then its extension $K(\alpha)$, where $\alpha\in\mathbb{C}$, is conjugation invariant if and only if $\bar{\alpha}\in K(\alpha)$.
Lemma 7.
Let $K\subseteq L$ be a radical extension of a conjugation invariant field $K$. Then there exist a chain $K=K_{0}\subseteq K_{1}\subseteq\cdots\subseteq K_{m}$ of simple radical extensions (of prime degree) with $L\subseteq K_{m}$, where all fields $K_{i}$ are conjugation invariant.
Proof.
We use induction with respect to the number $n$ of different simple radical extensions between $K$ and $L$. As already noted, we may assume that every simple radical extension is by a root of a binomial of prime degree. Asume that the Lemma is true for every radical extension of $K$ by such a chain of at most $n\geq 0$ of simple radical extensions. The base case, when $L=K$ (that is, $n=0$), is trivially true.
Let $L$ be a radical extension of $K$ by a chain of $n+1$ simple radical extensions (each by a root of a binomial of prime degree). Let $L^{\prime}\subset L$ be the last subextension in the chain. Then $K\subset L^{\prime}(\alpha)=L$, where $\alpha^{q}=a\in L^{\prime}$ for a prime number $q$. The field $L^{\prime}$ is a radical extension of $K$ by n1n-1 simple radical extensions. By the induction hypothesis, we have $L^{\prime}\subset K_{m^{\prime}}$, where $K=K_{0}\subset K_{1}\subset\cdots\subset K_{m^{\prime}}$ is a chain of simple radical extensions with all fields $K_{i}$ conjugation invariant and $K_{i}=K_{i-1}(\alpha_{i})$ with $\alpha_{i}^{q_{i}}=a_{i}\in K_{i-1}$ and $q_{i}$ a prime number for $i=1,\ldots,m^{\prime}$.
We construct a chain for $L$. We have that $L\subset K_{m^{\prime}}(\alpha)$. If $\bar{\alpha}\in K_{m^{\prime}}(\alpha)$, then $K_{m^{\prime}}(\alpha)$ is conjugation invariant and we put $K_{m^{\prime}+1}=K_{m^{\prime}}(\alpha)$. The chain
$K=K_{0}\subseteq\cdots\subseteq K_{m^{\prime}}\subseteq K_{m^{\prime}+1}.$
[…]
$K=K_{0}\subseteq\cdots\subseteq K_{m^{\prime}}\subseteq K_{m^{\prime}+1}\subseteq K_{m^{\prime}+2}\;.$
Clearly, we have that $L=L^{\prime}(\alpha)\subseteq K_{m^{\prime}}(\alpha)\subseteq K_{m^{\prime}}(\alpha,\bar{\alpha})=K_{m^{\prime}+2}$. If we have an equality of fields in the chain, we can simply remove repeating copies. The constructed chain satisfies the induction step. ∎
Remark 8.
This result is not formulated as such by Weber, but the idea of the construction is due to him. Weber [WW1, WW2] just says that he adjoins $\alpha$ and $\bar{\alpha}$ at the same time. He considers this clearly as one step, and only when needed he first adjoins $\rho$ and then $\alpha$. In particular, the first adjunction in a chain that makes $f(X)$ reducible is an extension $L\subset L(\alpha,\bar{\alpha})$ of conjugation invariant fields, which subsequently is studied in greater detail. The same holds for Dörrie []. The objection that sometimes has been raised to his proof, that not all extensions in the chain are conjugation invariant, is therefore not founded.
3. A proof of Kronecker’s Theorem.
In this section, we give a proof of Kronecker’s Theorem, which, in principle, allows us to treat both reducible and irreducible radicals using the same arguments. However, there is one point where a deeper knowledge of Galois theory for cyclic field extensions seems to be needed for such a unification. We comment on this after the proof, but we choose to give an elementary proof, where we consider the two cases (irreducible and irreducible) separately. For the second case we use an argument that more closely follows Weber’s ideas. We formulate Kronecker’s Theorem for polynomials with real coefficients in a conjugation invariant field.
Theorem 9 (Kronecker’s Theorem).
Let $K$ be a conjugation invariant number field. Suppose that an irreducible polynomial equation $f(X)\in K[X]$ of odd prime degree $p$ with real coefficients is solvable by radicals over $K$. Then exactly one root or all roots of the polynomial are real.
Proof.
We first remark that $f(X)$ has at least one real root, since it has real coefficients. As $f(X)$ is solvable, we may assume by Lemma 7 that there is a conjugation invariant chain $K=K_{0}\subseteq K_{1}\subseteq\cdots\subseteq K_{n}=L$ of simple radical extensions such that $K_{i}$ is an extension of $K_{i-1}$ by a root of binomial of prime degree for $i=1,\ldots,n$ and $K_{f}\subseteq L$. There exist an $i$ such that $f(X)$ is irreducible in $K_{i-1}$ but reducible in $K_{i}=K_{i-1}(\alpha_{i})$, where $\alpha_{i}^{q}=a_{i}\in K_{i-1}$ and $q$ is a prime number. For ease of notation we write $\alpha$ for $\alpha_{i}$ and $a$ for $a_{i}$. By Nagell’s Lemma 5 the degree $p$ of $f(X)$ divides $d=[K_{i}:K_{i-1}]$. We distinguish between two cases depending on whether $X^{q}-a$ is irreducible or $X^{q}-a$ is reducible over $K_{i-1}$.
Case I. If $X^{q}-a$ is irreducible over $K_{i-1}$, then $d=q$ is prime, which implies that $q=p$. We may assume that $K_{i-1}$ contains a primitive $p$-th root of unity $\varepsilon_{p}$, for otherwise we can replace $K_{i-1}$ by $K_{i-1}(\varepsilon_{p})$. Both polynomials $X^{p}-a$ and $f(X)$ remain irreducible over $K_{i-1}(\varepsilon_{p})$ by Nagell’s Lemma, as the degree $p$ is relatively prime to the degree of the minimal polynomial of $\varepsilon_{p}$ over $K_{i-1}$. In fact, this polynomial divides $X^{p-1}+X^{p-2}+\cdots+1$ (the minimal polynomial of $\varepsilon_{p}$ over $\mathbb{Q}$), so its degree is at most equal to p1p-1. Thus $K_{i}=K_{i-1}(\alpha)$ is a splitting field of the polynomial $X^{p}-a$ (see Lemma 4 (c)), which says that the extension $K_{i-1}\subset K_{i}$ is normal. Notice that in this case $K_{i}:K_{i-1}]=p$.
Since $f(X)$ is reducible over $K_{i}$ (and irreducible over $K_{i-1}$), the degrees of all irreducible factors of $f(X)$ are equal by Lemma 6. Thus the factors of $f(X)$ over $K_{i}$ are linear and all roots $\beta_{1},\dots,\beta_{p}$ of $f(X)$ belong to $K_{i}$. Hence, the splitting field $(K_{i-1})_{f}=K_{i-1}(\beta_{1},\beta_{2},\dots,\beta_{p})$ of $f(X)$ over $K_{i-1}$ is a subfield of $K_{i}$. For every $r=1,\ldots,p$, we have $K_{i-1}(\beta_{r}):K_{i-1}]=p$. It implies that $K_{i-1}(\beta_{r})=K_{i}$, since also $K_{i}$ has degree $p$ over $K_{i-1}$. Thus $K_{i}=(K_{i-1})_{f}=K_{i-1}(\beta_{r})$ for every $r=1,\ldots,p$.
Suppose that at least two roots, $\beta_{1}$ and $\beta_{2}$, are real. By Lemma 4 (c) the Galois group $G((K_{i-1})_{f}/K_{i-1})$ is cyclic of order $p$. Every nontrivial automorphism has order $p$ and since the group acts transitively on the roots $\beta_{r}$ (see Lemma 3), its elements are cycles of length $p$. We may choose as a generator the cycle $\sigma=(\beta_{1},\beta_{2},\dots,\beta_{p})$ starting with the real roots $\beta_{1},\beta_{2}$. Because the powers $\beta_{1}^{j}$ form a basis of $(K_{i-1})_{f}$ over $K_{i-1}$, we can write
(2) $\beta_{2}=c_{0}+c_{1}\beta_{1}+\ldots+c_{p-1}\beta_{1}^{p-1},$
with uniquely determined coefficients $c_{j}\in K_{i-1}$. Since $\beta_{1}$ and $\beta_{2}$ are real, complex conjugation gives
$\beta_{2}=\bar{c}_{0}+\bar{c}_{1}\beta_{1}+\ldots+\bar{c}_{p-1}\beta_{1}^{p-1},$
so $\bar{c}_{i}=c_{i}$ for $i=0,1,\ldots,p-1$. Thus all the coefficients $c_{i}$ are real.
Applying the automorphism $\sigma$ to equation (2), we get
$\beta_{3}=c_{0}+c_{1}\beta_{2}+\ldots+c_{p-1}\beta_{2}^{p-1},$
showing that $\beta_{3}$ is also real. By repeating this last argument, we find that all roots of $f(X)$ are real, if more than one root is real.
Case II. If $X^{q}-a$ is reducible, then $K_{i}=K_{i-1}(\varepsilon_{q})$ for a primitive $q$-th root of unity according to Lemma 4 (d). In order to simplify notations, we denote $\varepsilon_{q}$ by $\varepsilon$. Also in this case, the extension $K_{i-1}\subset K_{i}$ is normal as a splitting field of $X^{q}-1$ and $K_{i}:K_{i-1}]=d>1$.
Since the extension $K_{i-1}\subset K_{i}$ is normal and $f(X)$ is reducible over $K_{i}$ (and irreducible over $K_{i-1}$), we get as in Case I that all roots $\beta_{1},\dots,\beta_{p}$ of $f(X)$ belong to $K_{i}$. Hence the splitting field $(K_{i-1})_{f}=K_{i-1}(\beta_{1},\beta_{2},\dots,\beta_{p})$ of $f(X)$ over $K_{i-1}$ is a subfield of $K_{i}$. We work in the field $K_{i}$ instead of $(K_{i-1})_{f}$, which we did in Case I.
As we know, the polynomial $f(X)$ has a real root. Assume that $\beta_{1}$ is real. Since $K_{i}=K_{i-1}(\varepsilon)$, we can write
$\beta_{1}=c_{0}+c_{1}\varepsilon+\dots+c_{d-1}\varepsilon^{d-1}=\psi(\varepsilon_{q})\;,$
where $c_{r}\in K$ for $r=0,\dots,d-1$ and $\psi(X)=c_{0}+c_{1}X+\dots+c_{d-1}X^{d-1}$. Conjugating both sides in the last equalty, we get $\beta_{1}=\psi(\varepsilon)=\bar{\psi}(\bar{\varepsilon})=\bar{\psi}(\frac{1}{\varepsilon})$ (since $\varepsilon\bar{\varepsilon}=1)$. From Lemma 4 (d) it follows that every automorphism of $K_{i-1}(\varepsilon)$ maps $\varepsilon$ onto a power of this number. Thus the equality $\psi(\varepsilon)=\bar{\psi}(\bar{\varepsilon})=\bar{\psi}(\frac{1}{\varepsilon})$ still holds when $\varepsilon$ is replaced by another root of unity, which is the image of $\varepsilon$ by any automorphism of the field $K_{i-1}(\varepsilon)$. At the same time such an automorphism maps $\beta_{1}$ onto another root of the polynomial $f(X)$. Since this polynomial is irreducible over $K_{i-1}$, any root can be mapped at any other root (see Lemma 3), so the equality $\beta_{r}=\bar{\beta}_{r}$ is valid for all roots of $f(X)$, showing that they are real numbers. ∎
Remark 10.
In the proof above we have the situation that all roots $\beta_{1},\dots,\beta_{p}$ of an irreducible polynomial $f(X)\in K[X]$ generate subfields $K(\beta_{i})$ of an extension field $L\supseteq K$ of the same degree $p$ over $K$. In Case I, we conclude that all these fields are equal, since also $L$ has degree $p$ over $K$. In case II, it is not possible to claim the equality of all subfields $K(\beta_{i})$ using the same argument. But in both cases, the extension $K\subset L$ is normal and cyclic, so the Galois group is cyclic. A known property of finite cyclic groups says that for every divisor $d$ of the group order, there is exactly one subgroup whose order is equal to $d$. Using the Galois correspondence between the intermediate fields of a Galois extension $K\subset L$ (that is, fields $M$ such that $K\subseteq M\subseteq L$) and the subgroups of the Galois group $G(L/K)$, we get immediately that for every divisor $d$ of the degree of $L$ over $K$, there is exactly one field $M$ whose degree over $K$ is equal to $d$. If we know this in Case II (thus, assuming a little more knowledge of Galois theory, but not the correspondence between solvable groups and solvable extensions), we can conclude that the splitting field $(K_{i-1})_{f}$ is generated over $K_{i-1}$ by any of the roots of $f(X)$, and we can use the argument of Case I. Treating in this way both cases simultaneously results in a unified and short proof of Kronecker’s Theorem.
4. Weber’s proof.
In this Section, we return to the history of “elementary” proofs and discuss the errors that such proofs had until recently, and proposed corrections. At the end, we prove Kronecker’s Theorem using Weber’s method supplemented by the case of reducible radicals.
As our proof is inspired by the presentation of Nagell [Na], it follows the main lines of Weber’s proof. Given a chain $K=K_{0}\subset K_{1}\subset\cdots\subset K_{n}=L$ of simple radical extensions, the essential point is to study what happens in the first extension $K_{i-1}\subset K_{i}$ such that the solvable polynomial $f(X)$ of degree $p$ is irreducible in $K_{i-1}$ but reducible in $K_{i}$, where $K_{i}=K_{i-1}(\alpha)$ with $\alpha$ a root of the binomial $X^{q}-a$, $q$ prime. It is assumed that $K_{i-1}$ contains a $p$-th root of unity $\varepsilon_{p}$. Weber (and Epstein) [WW1, WW2] as well as Nagell [Na] claim that $q=p$, but this does not hold if $X^{q}-a$ is reducible over $K_{i-1}$ (in this case $\alpha$ is called a reducible radical). Weber and Dörrie give the following variant of Nagell’s Lemma 5 ([WW2, § 111, 5] and [, § 24, Satz IV]).
Proposition 11.
An irreducible polynomial $f(X)$ of prime degree $p$ can only become reducible through adjunction of a root of an [irreducible] equation $\varphi(X)=0$, whose degree is divisible by $p$.
We write the word irreducible in brackets, because it occurs in [], but it is left out in [WW2], although the proof in [WW2] uses the assumption that the equation $\varphi(X)=0$ is irreducible. The Proposition is in [WW2] applied to the equation $X^{q}=a$, with the conclusion that $q=p$. Irreducibility of this equation is not assumed, not even implicitly, as the next step argues that $a$ is not the $p$-th power of a number $\beta\in K_{i-1}$, because otherwise $\alpha=\varepsilon_{p}^{k}\beta$ would belong to $K_{i-1}$. This is clearly a circular argument.
Dörrie [] apparently noticed the error. To be able to use the correctly stated Proposition 11 he requires the radicals in the chain to be irreducible. The problem with reducible radicals was also commented on by Skopenkov [Sk, p. 183], who pointed out an error in the proof by Prasolov [Pr]. Prasolov follows Dörrie, but he does not include irreducibility in the definition of solvability by radicals; nevertheless, he states without justification that the polynomial $X^{q}=a$ is irreducible.
Dörrie does not notice that he might introduce reducible radicals when making the chain conjugation invariant:
Ferner wollen wir mit jedem adjungierten Radikal unsere Kette, das noch keine Zerlegung von $f(x)$ ermöglicht, auch gleich das komplex konjugierte Radikal adjungieren. Das ist vielleicht überflüssig, sicher aber nicht schädlich. [^6] Furthermore, with each adjoined radical of our chain, which not yet makes it possible to factorise $f(x)$, we will also adjoin at the same time the complex conjugate radical. That is maybe superfluous, but certainly not harmful.
But actually harm can be done, as the following example shows.
Example 12.
Let $\zeta$ be a primitive 7-th root of unity and let $f(X)$ be the minimal polynomial of the real number $\beta=\zeta+\bar{\zeta}$. One easily computes that
$f(X)=X^{3}+X^{2}-2X-1\;.$
In fact, $\beta$ generates the maximal real subfield of the cyclotomic field $\mathbb{Q}(\zeta)$ whose degree over $\mathbb{Q}$ is equal to 6 (the minimal polynomial of $\zeta$ is $(X^{7}-1)/(X-1)$). Thus its maximal real subfield has degree 3 over $\mathbb{Q}$. A similar construction is possible for any prime number $p$ such that $2p+1$ is also a prime number (Sophie Germain primes).
Let a chain of simple radical extensions start with $K_{0}=\mathbb{Q}\subset K_{1}=\mathbb{Q}(\alpha)$, where e.g. $\alpha=\zeta\sqrt[7]{11}$ and $\sqrt[7]{11}$ is the real root. Then $\alpha$ is an irreducible radical (the root of $X^{7}-11=0$) and the polynomial $f(X)$ of degree 3 is irreducible over $K_{1}$ by Nagell’s Lemma 5 as 3 does not divide the degree $K_{1}:\mathbb{Q}]=7$. But when we extend $K_{1}=\mathbb{Q}(\alpha)$ by the complex conjugate $\bar{\alpha}=\zeta^{6}\sqrt[7]{11}$, then $\bar{\alpha}$ is a reducible radical over $K_{1}$ (it satisfies the equation $X^{7}-11=0$ and the polynomial $X^{7}-11$ is reducible over $K_{1}$). Moreover, we have $\zeta\in K_{2}=\mathbb{Q}(\alpha,\bar{\alpha})$ as $\alpha/\bar{\alpha}=\zeta^{2}$ and $\zeta^{2}$ is a primitive $7$-th root of unity. Hence $\beta=\zeta+\bar{\zeta}\in\mathbb{Q}(\alpha,\bar{\alpha}))$ and $f(X)$ is reducible over this field. According to Weber and Dörrie and the construction in Lemma 7 we should in this case first adjoin $\alpha\bar{\alpha}=\sqrt[7]{121}$ and then $\alpha$, but $X^{7}-11$ is also reducible over $\mathbb{Q}(\sqrt[7]{121})$ as $\sqrt[7]{11}=(\sqrt[7]{121})^{4}/11$ is a root in $\mathbb{Q}(\sqrt[7]{121})$.
The polynomial $f(X)$ also provides an example for Weber’s and Nagell’s error, as it becomes reducible by the conjugation invariant extension $\mathbb{Q}\subset\mathbb{Q}(\zeta)$ of degree 6.
It is easy to solve the equation $f(X)=0$ using a computer algebra system. The three roots are
$\frac{\sqrt[3]{28+84\sqrt{-3}}}{6}+\frac{14}{3\sqrt[3]{28+84\sqrt{-3}}}+\frac{1}{3}\;,$
where one should always take the same values of the multiply occurring square and cubic roots. This answer avoids the ambiguities in Cardano’s formulas.
Weber’s and Dörrie’s argument can be rescued if the adjunction of a reducible radical making the polynomial $f(X)$ reducible is replaced by a chain of adjunctions of irreducible radicals, using Gauss’ result that the roots of unity are expressible by irreducible radicals [St, Ch. 21]. This is the approach chosen by Skopenkov [Sk] and Pan–Chen [PC]. Instead of the conjugation invariant chain provided by our Lemma 7 they construct a chain where in addition all prime degree binomials are irreducible ([PC, Thm. 5.1] and [Sk, Lemma 8.4.15]). The proof is by induction, using the fact that there exists for any prime $p\geq 3$ numbers $\beta_{0},\dots,\beta_{p-2}$ such that $\beta_{i}\in\mathbb{Z}[\varepsilon_{p-1}]$ with $\varepsilon_{p-1}$ a primitive (p1)(p-1)-th root of unity and $\varepsilon_{p}\in\mathbb{Q}(\varepsilon_{p-1},\beta_{0},\dots,\beta_{p-2})$ (see [St, (21.9)]). This procedure requires the adjunction of many roots of unity.
The complications caused by making all radicals irreducible can be avoided by using the same argument in the reducible case as in our proof of Theorem 9. That argument is an adaptation of Weber’s original arguments. He uses the assumption that $f(X)$ is a real polynomial in a different way than we do in the first case of the proof, but the arguments up to that point are quite similar (as our proof is inspired by Nagell’s [Na]). To prove Kronecker’s Theorem, using Weber’s method, we may thus assume that the polynomial $f(X)$ is irreducible over $K_{i-1}$, which we denote by $K$, that $K$ contains the $p$-th roots of unity, and that $f(X)$ splits in linear factors in $K_{i}=K(\alpha)=L$ (that is, all roots of $f(X)$ are in $L$). We also assume that both $K$ and $L$ are conjugation invariant. What remains to show is the following result, which we call lemma, as it is only the last part of Weber’s complete proof.
Lemma 13.
Let $f(X)\in K(X)$ be an irreducible polynomial of odd prime degree $p$ with real coefficients lying in a conjugation invariant field $K$ that contains a primitive primitive $p$-th root of unity $\varepsilon$, and let $K\subset L=K(\alpha)$, where $\alpha^{q}=a$, $a\in K$, be a normal and conjugation invariant extension such that $f(X)$ splits in linear factors over $L$. Then exactly one root or all roots of $f(X)$ are real.
Proof.
We distinguish two cases, depending on whether $X^{q}-a$ is irreducible or reducible.
Case I. If $X^{q}-a$ is irreducible over $K$, then $q=p$. As $\varepsilon\in K$, the extension $K\subset L$ is normal as a splitting field of $X^{p}-a$. All roots $\beta_{1},\dots,\beta_{p}$ of the polynomial $f(X)$ are in $L$. Since $f(X)$ has real coefficients and odd degree, at least one of its roots is real. We assume that it is $\beta_{1}$. All roots can be expressed as linear combinations of $1,\alpha,\dots,\alpha^{p-1}$. Therefore we have
(3) $\beta_{1}=c_{0}+c_{1}\alpha+\dots+c_{p-1}\alpha^{p-1}\;,$
where $c_{i}\in K$. Now we distinguish two cases depending on whether $\alpha\in\mathbb{R}$ or $\alpha\notin\mathbb{R}$.
Case Ia. If $\alpha\in\mathbb{R}$, then the complex conjugation of equation (3) gives
$\beta_{1}=\bar{c}_{0}+\bar{c}_{1}\alpha+\dots+\bar{c}_{p-1}\alpha^{p-1}\;,$
and the $\bar{c}_{i}$ also lie in $K$. By uniqueness of the representation $\bar{c}_{i}=c_{i}$, or in other words, all $c_{i}$ are real. As the automorphisms of the field $L$ over $K$ map $\alpha$ onto $\varepsilon^{r}\alpha$ for $r=0,\dots,p-1$, the other roots of $f(X)$ are
$\beta_{r}=c_{0}+c_{1}\varepsilon^{r}\alpha+\dots+c_{p-1}\varepsilon^{r(p-1)}\alpha^{p-1}\;.$
As $\bar{\varepsilon}^{r}=\varepsilon^{p-r}$, we find that $\bar{\beta}_{r}=\beta_{p-r}$, so the other roots come in complex conjugate pairs and in this case $\beta_{1}$ is the only real root.
Case Ib. If $\alpha\notin\mathbb{R}$, then consider $\rho=\alpha\bar{\alpha}\in L$. We have $K\subseteq K(\rho)\subseteq L$ and since $[L:K]=p$, we have either $K(\rho)=L=K(\alpha)$ or $K(\rho)=K$. If $K(\rho)=K(\alpha)$, we can replace $\alpha$ by the real number $\rho$ and we are in Case Ia, since $\rho^{p}=(\alpha\bar{\alpha})^{p}=a\bar{a}$ is real (in fact, it follows that $a$ is real in this case). If $K(\rho)=K$, then $\rho=\alpha\bar{\alpha}\in K$. The equality $\rho=\alpha\bar{\alpha}$ gives $\bar{\alpha}=\frac{\rho}{\alpha}$ and $\bar{\alpha}^{r}=\frac{\rho^{r}}{\alpha^{r}}=\frac{\rho^{r}}{a}\alpha^{p-r}$, so in this case complex conjugation of equation (3) for the real root $\beta_{1}$ gives $c_{0}=\bar{c}_{0}$ and $\bar{c}_{r}=\frac{a}{\rho^{r}}c_{p-r}$. For the other roots $\beta_{s}$, we find $\bar{\varepsilon}^{rs}\bar{\alpha}^{r}=\frac{\rho^{r}}{a}\varepsilon^{(p-r)s}\alpha^{p-r}$, which implies $\bar{\beta}_{s}=\beta_{s}$. Therefore all roots are real. This explicit computation makes the argument similar to the Case Ia.
Weber argues that $\beta_{0}=\psi(\alpha)=\bar{\psi}(\bar{\alpha})=\bar{\psi}(\frac{\rho}{\alpha})$, where $\psi(X)=c_{0}+c_{1}X+\dots+c_{p-1}X^{p-1}$ and as $\bar{\alpha}_{s}=\frac{\rho}{\alpha_{s}}$ for any root of $X^{p}-a$, the equality $\psi(\alpha)=\bar{\psi}(\bar{\alpha})$ still holds when $\alpha$ is replaced by another root when we apply the automorphism mapping $\beta_{1}$ onto this root (that is, we map $\alpha$ onto $\varepsilon^{s}\alpha$).
Case II. If $X^{q}-a$ is reducible over $K$, then $\alpha^{q}=a_{0}^{q}$ for some $a_{0}\in K$ by Lemma 4 (b). Therefore $\alpha=\varepsilon_{q}a_{0}$ for some primitive $q$-th root of unity and $K(\alpha)=K(\varepsilon_{q})$. This situation is exactly the one treated in Case II in the proof of Kronecker’s Theorem in Section 3; one has to read $K$ for $K_{i-1}$ and $L$ for $K_{i}$. ∎
Acknowledgements.
We express our gratitude to the referee for all remarks and comments, which contributed to a significant improvement of the paper.